If three rectangular holes of 1.5m x 0.3m x 200mm are placed on
the wall,
1. How many cubic meters of concrete needed to build the wall?
2. What is the total volume of the rectangular solid holes?
1
Expert's answer
2020-10-11T17:37:16-0400

"1\hspace{0.1cm} \textsf{mm} = 10^{-3}\hspace{0.1cm} \textsf{m}\\


200\hspace{0.1cm}\textsf{mm} = 200 \times 10^{-3} \hspace{0.1cm}\textsf{m}= 0.2 \hspace{0.1cm}\textsf{m}\\

\therefore\textsf{The total volume of the}\\\textsf{rectangular solid holes is}\\ \textsf{length}\times \textsf{breadth}\times \textsf{height} =\\1.5\hspace{0.1cm}\textsf{m} \times 0.3\hspace{0.1cm}\textsf{m} \times 0.2 \hspace{0.1cm}\textsf{m} = 0.09 \hspace{0.1cm}m^3\\



\textsf{And the cubic meters of concrete}\\\textsf{needed to build the wall is}\\ 3 \times 0.09\hspace{0.1cm}m^3 = 0.27\hspace{0.1cm}m^3"